OpenStudy (anonymous):

The ratio of two side lengths of the triangle is given. Solve for the variable, then find the length of CB . AB:CB is 1:3 please help

OpenStudy (anonymous):



OpenStudy (amistre64):

ab/cb = 1/3 cb = 3ab

OpenStudy (anonymous):

sorry for wat i said earlier. was having problems with a girl and can you explain what tht means

OpenStudy (amistre64):

the ratio they give you suggest a fraction\[\frac{ab}{cb}=\frac13\]rewriting this we get:\[3ab=cb\] input the values of ab and cb that are in the problem 3x = 2x+3 and solve for x

OpenStudy (amistre64):

once you know x, you can then determine the specific values for ab and cb

OpenStudy (anonymous):

im rlly slow st this....... how would i find x

OpenStudy (amistre64):

this is as basic as algebra gets ... 3x = 2x+3 what are some ideas you have that might be useful in solving it?

OpenStudy (anonymous):

2x+3/3x this is as basic as i suck im sorry

OpenStudy (anonymous):

? after 3x

OpenStudy (amistre64):

:) lets start by getting the xs to one side .... id suggest subtract 2x from each side 3x = 2x+3 -2x -2x ------------ what do we get here?

OpenStudy (anonymous):

1x=3???

OpenStudy (amistre64):

correct so, since x = 3 and cb = 2x+3 or cb = 3x what is the value of cb?

OpenStudy (anonymous):

im so fluttering stupid... 9 thankyo u

OpenStudy (anonymous):

i think you can guess wat fluttering was supposed to be lol